OpenStudy (anonymous):

X2 - 3x - 54 show all working please ^_^

OpenStudy (uri):

use the quadratic formula mate,

OpenStudy (uri):

\[x^2-3x-54\]

OpenStudy (uri):

a = 1 ; b = -3 ; b = -54 and plug it into the quad. formula.

OpenStudy (uri):

c=-54**

OpenStudy (anonymous):

Im gonna try ............Thank you =)

OpenStudy (uri):

:)

OpenStudy (shubhamsrg):

or simply, -54 = -9 * 6 and -3 = -9+6

OpenStudy (anonymous):

Thank you =)